OpenStudy (anonymous):

Verify the identity. - tan^2(x) + sec^2(x) = 1

OpenStudy (anonymous):

sin^(x)+cos^2(x)=1 ] divide both sides by cos^(x) you will get the required answer...

OpenStudy (anonymous):

if it is not clear let me know to prove it for u in details...

OpenStudy (mertsj):

tan^2x= sec^2x-1

OpenStudy (mertsj):

\[-(\sec ^2x-1)+\sec ^2x=1\]

OpenStudy (mertsj):

\[-\sec ^2x+1+\sec ^2x=1\]

OpenStudy (mertsj):

\[-\sec ^2x+\sec ^2x+1=1\]

OpenStudy (mertsj):

\[0+1=1\]

OpenStudy (mertsj):

\[1=1\]

OpenStudy (anonymous):

@Mertsj are u sure that's how?

OpenStudy (mertsj):

There are multiple ways. That is one of them. You could also do what mathsmind said or you could replace sec^2x with 1+tan^2x

OpenStudy (anonymous):

yes both ways are correct...